Solve the inequality

x>√(x+2)

IxI / Ix+1I -1 > 1


1
Expert's answer
2021-09-01T11:28:03-0400

1.


"x>\sqrt{x+2}"

We know that

"\begin{matrix}
 x+2\geq0 \\ \\
 \sqrt{x+2}\geq0
\end{matrix}"

Then we have "x>0"


"x^2>x+2"

"x^2-x-2>0"

"(x+1)(x-2)>0"

"x<-1\ or \ x>2"

Since "x>0," then we take "x>2."


Answer: "x>2."


"x\in(2, \infin)."


2.


"\dfrac{|x|}{|x+1|-1} > 1"

"|x+1|-1\not=0=>x\not=0, x\not=1"


"\begin{cases}
 |x+1|-1>0 \\
 |x|>|x+1|-1
\end{cases}"

"\begin{cases}
 |x+1|>1 \\
 |x|+1>|x+1|
\end{cases}"

"x\leq-1, x\not=-2"

"\begin{cases}
 -x-1>1 \\
 -x+1>-x-1
\end{cases}"

"\begin{cases}
x<-2 \\
 2>0
\end{cases}"

"x\in(-\infin, -2)"


"-1\leq x<0"

"\begin{cases}
 x+1>1 \\
 -x+1>x+1
\end{cases}"

"\begin{cases}
x>0 \\
 2x<0
\end{cases}"

No solution



"x>0"

"\begin{cases}
 x+1>1 \\
 x+1>x+1
\end{cases}"

"\begin{cases}
x>0 \\
 1>1
\end{cases}"

No solution


Answer: "x<-2"


"x\in(-\infin, -2)"